OpenStudy (kuvyojhmoob):

A single electron has a charge of 3.7×10−15 Z . Using zeets and a fictitious unit of mass called the wiggle, w, the charge-to-mass ratio of an electron is 3.9×106 Z/w . What is the mass of an electron in wiggles?

OpenStudy (anonymous):

assuming Z stands for charge $$\large 3.7\times 10^{-15}Z \cdot \frac {1 w }{3.9 \times 10^6 Z }= ~?$$

OpenStudy (kuvyojhmoob):

w=9.487*10^-22